1.     Determine the pH of a buffer prepared by adding 15 mL of 0.25 M acetic acid to 35 mL of 0.17 M sodium acetate. Given Ka for acetic acid = 1.7 x 10-5.


1
Expert's answer
2022-06-14T04:30:04-0400

Solution:

acetic acid = CH3COOH

sodium acetate = CH3COONa


Total volume of solution = Volume of CH3COOH + Volume of CH3COONa

Total volume of solutio = 15 mL + 35 mL = 50 mL


Calculate new concentrations considering dilution:

[CH3COOH] = (0.25 M) × (15 mL / 50 mL) = 0.075 M

[CH3COONa] = (0.17 M) × (35 mL / 50 mL) = 0.119 M


The Henderson-Hasselbalch equation is commonly used to calculate the pH of a buffer solution from the concentration of the buffer components:



where pKa is the acid dissociation constant and [base] and [acid] are the base and acid concentrations, respectively.

In this case, the base is CH3COONa, and the acid is CH3COOH.

Therefore,

pH = −log(1.7 × 10−5) + log(0.119 M / 0.075 M) = 4.97

pH = 4.97


Answer: pH = 4.97